Problem 6.14 Humidity and Condensation Your answer is partially correct. Try again. Air at 70.0% relative humidity is cooled isobarically at 1 atm absolute from 85.0°C to 35.0°C. a. Estimate the dew point and the degrees of superheat of the air at 85.0°C. Dew point: 76.2 loc 吖: loc 8.8 Degrees of superheat:
Answer #1

Given: Relative humidity, hr =70%=0.70

Initial temperature of air, T1=85 0C=358.15 K

Final temperature of air, T2=35 0C=308.15 K

Pressure, P=1 atm=760 mm Hg

(a) Dew point and Degree of superheat

Obtain the vapor pressure at 85 0C from the Vapor Pressure of Water Table.

At 800C, vapor pressure, p*=355.1 mm Hg and at 900C, p*=525.76 mmHg. Interpolate the values

Thus vapor pressure at 85 0C, p*(85 0C)=440.43 mm Hg

Mole fraction of water, yH2O, is

hrp(85°C no =

where hr is the Relative humidity

P is the Pressure

\Rightarrow y_{H_{2}O}=\frac{0.7*440.43\ mmHg}{760\ mmHg}

ун,0-0.4056 mol H2O/ mol

Dew point temperature, Tdp, is

y_{H_{2}O}P=p^{*}(T_{dp})

where p*(Tdp) is the Vapor pressure at dew point temperature

\Rightarrow p^{*}(T_{dp})=0.4056\ mol\ H_{2}O/mol*760\ mmHg

\Rightarrow p^{*}(T_{dp})=308.256\ mmHg

Refer Vapor Pressure of Water Table and obtain the temperature corresponding to 308.256 mmHg.

At 700C, p*=233.7 mmHg and at 800C, p*=355.1 mmHg. Interpolate the values.

The vapor pressure is 308.256 mmHg at a temperature of 76.140C

Dew point temperature, Tdp=76.140C

Degree of superheat=T1-Tdp=(85-76.14)0C

Degree of Superheat=8.860C

(b) To calculate water condensed per m3 of feed gas

Basis: 1 m3 feed gas

Number of moles of gas in feed,n, is obtained using the gas law

P'V'=n'RT'\ and\ PV=nRT_{1}

where P', V', n' and T' is the Pressure, Volume, Moles and Temperature at standard conditions

V is the Volume of feed gas

\Rightarrow n=\frac{VT'n'}{T_{1}V'}

\Rightarrow n=\frac{1\ m^{3}feed\ gas*273.15\ K*1\ mol}{358.15\ K*0.0224\ m^{3}(STP)}

\boldsymbol{n=34.029\ mol}

At saturation conditions, the mole fraction of water, y, is

y=\frac{p^{*}(35^{\circ}C)}{P}

where p*(350C) is the Vapor pressure of water at 350C, p*(350C)=42.175 mmHg

\Rightarrow y=\frac{42.175\ mmHg}{760\ mmHg}

\boldsymbol{y=0.056\ mol\ H_{2}O/mol}

Dry air balance is

(1-y_{H_{2}O})n=n_{1}(1-y)

(1-0.4056)n=n_{1}(1-y)

\Rightarrow 0.5944\ mol\ dry\ air/mol*34.029\ mol=n_{1}(1-0.056\ mol\ H_{2}O/mol)

\Rightarrow 20.2268=0.944n_{1}

\boldsymbol{n_{1}=21.427\ mol}

Total mole balance is

34.029\ mol=n_{1}+n_{2}

where n2 is the Moles of water condensed

\Rightarrow 34.029\ mol=21.427\ mol+n_{2}

\mathbf{n_{2}=12.602\ mol\ H_{2}O/m^{3}}

Thus moles of water condensed is n2=12.602 mol H2O/m3 feed gas

(c) Pressure at which mist forms is obtained using Raoult's law

y_{H_{2}O}P=p^{*}(85^{\circ}C)

\Rightarrow P=\frac{p^{*}(85^{\circ}C)}{y_{H_{2}O}}

\Rightarrow P=\frac{440.43\ mmHg}{0.4056\ mol\ H_{2}O/mol}

\Rightarrow P=1085.87\ mmHg

\mathbf{P=1.43\ atm}

Thus the pressure at which mist forms is P=1.43 atm.